Using a table of thermochemical data, calculate Enthalpy for the reaction.

CS2(g)+2H2O(l)=CO2(g)+2H2S(g)


1
Expert's answer
2021-06-14T11:36:01-0400

full question

Calculate the enthalpy of the following chemical reaction: [6]

CS2(ℓ) + 3O2(g) → CO2(g) + 2SO2(g)

Given:

C(s) + O2(g) → CO2(g) ΔH = -393.5 kJ/mol

S(s) + O2(g) → SO2(g) ΔH = -296.8 kJ/mol

C(s) + 2S(s) → CS2(ℓ) ΔH = +87.9 kJ/mol

Answer:-

"\Delta{H}_{reaction}=\sum{\Delta{H_{products}-\sum\Delta{H}_{reactants}}}=-393.5+2(-296.8)-87.9=-1075kJ/mol"


Answer: -1075 kJ/mol